please help to solve the following:
TR=110Q-Q2
TC=1/2Q2+50Q+30
Determine the level at which the firm maximizes its total profit.

Answer by stanbon(75887)

solve the following:
TR=110Q-Q2
TC=1/2Q2+50Q+30
Determine the level at which the firm maximizes its total profit.
----
PR = TR-TC
PR = -(3/2)Q^2 + 60Q - 30
----
max occurs when Q = -b/2a = -60/(2(-3/2)) = 20
---
Maximized profit when Q == 20
--------------------------------
Cheers,
Stan H.